OpenStudy (anonymous):

Find the indicated term. a1 = 9, r = e^−x, n = 7 find a7

OpenStudy (anonymous):

\[a_n=ar^{n-1}\] \[a_7=ar^6\]

OpenStudy (anonymous):

yeah I get that, but what do I do with the e^-x? am I supposed to figure out what e^-x is or? I tried solving and writing in the actual number but it was wrong.

OpenStudy (anonymous):

\[a_7=9(e^{-x})^6\] we can only write in terms of x

OpenStudy (anonymous):

thank you!!!! I feel like an idiot! it was that simple!!! gah! thank you!

OpenStudy (anonymous):

medal